OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = x2 y = x + 2

OpenStudy (anonymous):

set them equal to each other:x^2=x+2 x={-1,2} Them subs these values into either one of the equations: y=(-1)^2=1-->(-1,1) y=(2)^2=4---->(2,4)